OpenStudy (anonymous):

the distance around the inside of a shopping mall is is 12/16 mi aaron walked around the mall 3 times how far did he walk

OpenStudy (anonymous):

12/16 + 12/16 + 12/16 = 36/16 = 2 + 4/16 = 2 + 1/4